# Performance Comparison of HTTP/3 and HTTP/2: Proxy vs. Non-Proxy Environments

### v1

@@ Abstract.
This paper provides a systematic evaluation of the performance of QUIC/HTTP3(H3) and TCP/HTTP2(H2) protocols in proxy-enhanced environments. By leveraging features such as UDP-based flow-controlled streams, integrated TLS, multiplexed connections, and connection migration, H3 promises enhanced web communication. Despite extensive research, the impact of proxy integration and connection migration remains underexplored. This study addresses this gap by evaluating these protocols across various scenarios in noisy networks and proxy setups. Our findings reveal that H3 excels under high loss and latency conditions, significantly benefiting from its connection migration and multiplexing features. H3’s connection migration remains robust, maintaining stable performance even in proxy-enhanced environments, ensuring seamless network transitions. The proxy has a more neutral impact on H3, while it significantly enhances H2 performance, especially when using BBR. Any improvements observed in H3 under a proxy are minor and do not fundamentally alter H3’s performance as they do for H2. Importantly, while H2 with the right congestion control algorithm (CCA) can achieve performance comparable to H3, H3’s performance is more robust, as it is less impacted by network conditions, proxy settings, and CCA variations.
@@ 1. Introduction
QUIC introduces several innovative features that distinguish it from previous protocols. It enables faster connection establishment and supports multiplexed connections without head-of-line (HoL) blocking, a substantial limitation in TCP. Furthermore, QUIC uniquely supports connection migration, which allows ongoing sessions to continue even when the client’s network endpoint changes (e.g., when switching from Wi-Fi to cellular networks) (Iyengar and Thomson, 2022). HTTP/3 builds on QUIC by incorporating these advancements to streamline data transmission processes, enhancing web page load times and responsiveness through smarter packet processing and prioritization of requests (Bishop, 2022).
Major companies such as Google (chr, 2013), Akamai (Bishop, 2021), Meta (Joras and Chi, 2022), Microsoft (Banks, 2024), and Cloudflare (Ghedini and Lalkaka, 2019) have adopted QUIC commercially, indicating its wide adoption and significant impact on the global web infrastructure. As of April 1, 2024, 71.4% of desktop browsers and 70.8% of mobile browsers support H3 among the top 1000 websites (htt, 2024), underscoring its growing relevance.
@@
First-of-its-Kind Analysis of Connection Migration under Proxy Conditions: To the best of our knowledge, this paper is the first to evaluate QUIC’s connection migration capabilities under proxy conditions, including scenarios involving IP address changes and network impairments. Our results demonstrate H3’s ability to maintain low request times and ensure seamless user experience during network transitions, highlighting its suitability for mobile and dynamic networking environments.
@@ 2. Related Work
Performance of QUIC Extensive research has explored QUIC’s enhancements over traditional protocols, emphasizing its ability to significantly improve web browsing experiences. Early investigations by Carlucci et al. (Carlucci et al., 2015) demonstrated QUIC’s superior performance over TCP and SPDY (Wang et al., 2014), particularly in terms of connection setup times and robustness to packet loss. Cunha et al. (Da Cunha et al., 2023) further highlighted the conditions in which QUIC excels and when it underperforms. QUIC’s evolution has led to efforts investigating interoperability across implementations (Jaeger et al., 2023; Marx et al., 2020; Seemann and Iyengar, 2020). Long-term studies by Kakhki et al. (Kakhki et al., 2017) and Piraux et al. (Piraux et al., 2018) traced QUIC’s progression over time. Additionally, Rüth et al. (Rüth et al., 2018) examined QUIC deployment and usage in its early stages, while QScanner (Zirngibl et al., 2021) analyzed early QUIC deployments.
@@
Features of QUIC Research on H2 and H3, such as the studies by Marx et al. (Marx et al., 2019) and Sander et al. (Sander et al., 2022), has provided in-depth analysis of how resource multiplexing and prioritization affect web performance. These studies highlight that while H3’s use of QUIC reduces HoL blocking compared to H2, real-world gains depend heavily on effective prioritization strategies. Other research has extended QUIC to multipath QUIC (MPQUIC) (Coninck and Bonaventure, 2017; Ingenzi et al., [n. d.]), and new packet schedulers for MPQUIC (do Prado Paiva et al., 2023; Rabitsch et al., 2018; Wang et al., 2019) and upper-layer solutions (Zheng et al., 2021) have followed. MultipathTester (Coninck and Bonaventure, 2019) compares the performance of MPTCP with MPQUIC on iOS. Several studies have focused on QUIC congestion control (Engelbart and Ott, 2021; Haile et al., 2022; Mishra et al., 2022), while others examine security and privacy (Kosek et al., 2022c; Lychev et al., 2015; Nawrocki et al., 2021, 2022; Soni and Rajput, 2020). Experimental protocols inspired by QUIC, like TCPLS (Rochet et al., 2021), DCQUIC (Tan et al., 2021), and MCQUIC (Navarre et al., 2023), further show its potential.
@@
QUIC with Proxy Exploring QUIC in proxy architectures is relatively sparse, but some studies have started to bridge this gap. Kosek et al. (Kosek et al., 2023) introduced Secure Middlebox-Assisted QUIC (SMAQ), enabling middleboxes in QUIC connections while preserving security, and demonstrated performance improvements in satellite networks. Another study by Kosek et al. (Kosek et al., 2022a) evaluated QUIC and HTTP/3 with and without performance enhancing proxies (PEPs) in satellite networks, showing significant performance gains, especially in high-latency conditions. Kühlewind et al. (Kühlewind et al., 2021) assessed the QUIC-based MASQUE proxying protocol for tunneling UDP and IP traffic, revealing benefits like local recovery in lossy links despite introducing overhead. However, these studies focus on different aspects of QUIC and proxies. Our work uniquely compares H2 and H3 in proxy-enhanced environments, with a particular emphasis on QUIC’s connection migration feature, evaluating the performance impacts of proxy and connection migration under various network conditions.
@@
Single Stream Request Proxygen (H3) Curl (H2, H3) Chrome (H2, H3) Wget (H2) Quiche (H3) aioquic (H3) Nginx Bandwidth: 10 Mbps Loss: 0, 2, 4(%) Delay: 0, 25, 50(ms) Small Files: 0KB, 10KB, 100KB Medium Files: 500KB, 1MB, 5MB Large Files: 10MB Total Request Time
@@
Web page Loading Chrome (H2, H3) Nginx Bandwidth: 10 Mbps Loss: 0, 2, 4, 8, 12(%) Delay: 0, 25, 50(ms) 18 popular local host websites Speed Index
3. Repeatable Experimental Setup
We use a hardware network test bed that is remotely accessible for networking researchers and educators (Wiseman et al., 2008). The client and proxy machines run Ubuntu 20.04.6 LTS on Intel(R) Xeon(R) E5520 CPUs with 16GB RAM, while the server runs Ubuntu 22.04.4 LTS on an Intel(R) Core(TM) i3-3240 CPU with 16GB RAM. These machines are interconnected through a router via a 1 Gbps Ethernet network. The ’tc netem’ utility (tc, 2024) is used for bandwidth control, packet loss, and delay manipulation. In all experiments, the bandwidth is set to 10 Mbps, with specific loss and delay parameters detailed in each experiment’s setup (see Table 1). We use a fixed bandwidth to maintain controlled conditions, allowing us to isolate the effects of proxy environments on the performance of QUIC and HTTP/2. These bandwidth limitations, combined with varied loss and delay parameters on both uplink and downlink, represent realistic edge-case network conditions that induce unique protocol behaviors for root cause analysis. The study by Zhang et al. (Zhang et al., 2024) has already explored performance under varying bandwidth, while both our work and studies like (Yu and Benson, 2021) use fixed bandwidth with varied loss and delay to focus specifically on how network impairments affect protocol performance, without introducing the additional complexity of fluctuating bandwidth.
@@ 3. Repeatable Experimental Setup
In each experiment, we conducted 10 runs. We use Total Time as the metric in single-stream, multi-stream, and connection migration experiments, measuring from the client’s first request packet to the server until the client closes the connection. For web page loading, we use the Speed Index (spe, 2024), a metric that measures how quickly content appears on the user’s screen.
@@ 3.1. Baseline Experiment Setup
This experiment establishes the baseline for evaluating the performance of H2 and H3 in single-stream request, multi-stream request, connection migration, and web page loading. We employ Cubic as the CCA on the client side, and New Reno on the server side for both QUIC and TCP. Cubic is the default CCA for most of the H3 clients and the default CCA on Linux. For QUIC, the New Reno is the only CCA that Nginx supports. Both H2 and H3 use TLS 1.3. Our experimental setup is detailed in Table 1, with the topology illustrated in Figure 1.
Figure 1. Baseline Experiment Topology
@@
In the connection migration experiment, We extended Cloudflare’s Rust-based quiche client to support connection migration. Specifically, we took inspiration from the design of Andriod’s Connection Manager (Con, 2024) to detect changed interfaces/IP addresses, and initiate the connection migration process as described in (Iyengar and Thomson, 2022). For H2, we employed wget with and without Range Requests (ran, 2024), along with curl tests without Range Requests. Range requests in H2 facilitate fetching a range of data. This allows us to explore H2’s ability to pause and resume downloading in the change in connection scenario. The experiment simulates IP address changes after downloading over 2MB of a 5MB test file to mimic real-world network transitions, evaluating how each protocol maintains connection stability and download continuity during these transitions.
@@ 3.2. Proxy Experiment Setup
This experiment assesses the impact of introducing an Envoy reverse proxy between the server and client, with loss and delay applied between the proxy and the client. The reverse proxy breaks the connection between the client and the server and establishes a new connection with the server. We use BBR as the CCA for both QUIC and TCP since it is the default CCA for QUIC on Envoy. We disabled the proxy cache and retry features. The connection bandwidth is uniformly set at 10 Mbps across all links. We chose Envoy for its robust, production-grade capabilities and widespread adoption by major companies like Google, Microsoft, and Amazon. The network topology positions the Envoy between the client and server, as detailed in the accompanying Figure 3.
Figure 3. Proxy Experiment Topology
@@
Figure 4. Single Stream Results. The times represent the mean values over 10 runs.
@@ 4. Experimental Evaluation
In this section, we present the results from our experiments and our analysis of the results.
4.1. Single Stream Experiment Evaluation
Figure 4 presents the results of our single stream experiments across various network conditions, highlighting the performance differences between H2 and H3.
@@ 4.1. Single Stream Experiment Evaluation
Moderate Network Impairments In scenarios involving moderate impairments (10 Mbps bandwidth, 0% packet loss, and 50 ms delay), both H2 and H3 demonstrated comparable performance without the use of a proxy (Figure 4(b)). However, The introduction of a proxy caused slight variations in performance for H3 clients, such as curl and quiche, but overall, H3 maintained consistent performance with minimal impact from the proxy. In contrast, H2’s performance remained stable across different clients, suggesting that H2 implementations are more mature and better optimized for proxy environments. This indicates that further optimization is needed for some of the H3 clients to handle proxy-related challenges effectively.
In scenarios with loss only (10 Mbps bandwidth, 2-4% packet loss, 0 ms delay), both H2 and H3 exhibited comparable performance, with minimal impact from proxy use. This suggests that both protocols efficiently handle packet loss under these specific conditions.
@@
Connection Establishment The QUIC connection starts with a handshake that merges the TCP 3-way handshake and the TLS 1.3 handshake into a single three-packet exchange. This process eliminates a full Round-Trip Time (RTT) in the QUIC startup phase, offering significant performance improvements for short sessions. While both protocols show similar connection establishment times under ideal network conditions, H3’s advantages become evident under impaired conditions (Figure 4(c)). The quicker connection establishment does not significantly impact performance in optimal conditions, but it becomes crucial in severely impaired networks.
@@
Impact of CCAs In proxy scenarios, the use of BBR has shown to maintain high throughput and low latency by dynamically adjusting the congestion window based on real-time bandwidth and RTT estimations. This proactive congestion management is more effective than loss-based algorithms like Cubic, which only reduce the congestion window upon detecting packet loss. While BBR improves H2’s performance significantly in proxy-enhanced environments, this improvement is highly dependent on the correct CCA choice. In contrast, H3’s performance remains robust across different network conditions and CCAs, highlighting its superior consistency without relying on external configurations like BBR.
@@
Figure 5. Multi Stream Results. Each point represents the highest Total Request Time among 20 test files for each run. A total of 10 runs were conducted for each loss level.

### v2

@@ Abstract.
This paper systematically evaluates the performance of QUIC/HTTP3 (H3) and TCP/HTTP2 (H2) protocols in proxy-enhanced environments. H3 leverages features such as UDP-based flow-controlled streams, integrated TLS, multiplexed connections, and connection migration, offering the potential for improved web communication. Despite extensive research, the impact of proxy integration and connection migration remains underexplored. This study addresses this gap by evaluating H3 and H2 across various scenarios, particularly in noisy networks and proxy setups. Our findings show that H3 excels under high loss and high latency conditions, significantly benefiting from its connection migration and multiplexing features, with improvements of up to 88.36% under high-loss and high-latency conditions, and 81.5% under extreme loss conditions, respectively. H3’s connection migration remains robust, maintaining stable performance even in proxy-enhanced environments, ensuring seamless network transitions. The proxy has a more neutral impact on H3, while it significantly enhances H2 performance, particularly when paired with BBR, resulting in a 90% improvement in the single-stream file download experiment under severe network impairments. Any improvements observed in H3 under a proxy are minor and do not fundamentally alter H3’s performance as they do for H2. Importantly, while H2 with the right congestion control algorithm (CCA) can achieve performance comparable to H3, H3’s performance is more robust, as it is less impacted by network conditions, proxy settings, and CCA variations.
@@ 1. Introduction
QUIC introduces several innovative features that distinguish it from other previous protocols. It enables faster connection establishment and supports multiplexed connections without head-of-line (HoL) blocking, a substantial limitation in TCP. Furthermore, QUIC uniquely supports connection migration, which allows ongoing sessions to continue even when the client’s network endpoint changes (e.g., when switching from Wi-Fi to cellular networks) (Iyengar and Thomson, 2022). HTTP/3 builds on QUIC by incorporating these advancements to streamline data transmission processes, enhancing web page load times and responsiveness through smarter packet processing and prioritization of requests (Bishop, 2022).
Major companies such as Google (chr, 2013), Akamai (Bishop, 2021), Meta (Joras and Chi, 2022), Microsoft (Banks, 2024), and Cloudflare (Ghedini and Lalkaka, 2019) have adopted QUIC commercially, indicating its wide adoption and significant impact on the global web infrastructure. As of April 1, 2024, 71.4% of desktop browsers and 70.8% of mobile browsers support HTTP/3 among the top 1000 websites (htt, 2024), underscoring its growing relevance.
@@
First-of-its-Kind Analysis of Connection Migration under Proxy Conditions: To the best of our knowledge, this paper is the first to evaluate QUIC’s connection migration capabilities under proxy conditions, including scenarios involving IP address changes and network impairments. Our results demonstrate H3’s ability to maintain low request time and ensure seamless user experience during network transitions, highlighting its suitability for mobile and dynamic networking environments.
@@ 2. Related Work
Performance of QUIC Extensive research has explored QUIC’s enhancements over traditional protocols, emphasizing its ability to significantly improve web browsing experiences. Early investigations by Carlucci et al. (Carlucci et al., 2015) demonstrated QUIC’s superior performance over TCP and SPDY (Wang et al., 2014), particularly in terms of connection setup time and robustness to packet loss. Cunha et al. (Da Cunha et al., 2023) further highlighted the conditions in which QUIC excels and when it underperforms. QUIC’s evolution has led to efforts investigating interoperability across implementations (Jaeger et al., 2023; Marx et al., 2020; Seemann and Iyengar, 2020). Long-term studies by Kakhki et al. (Kakhki et al., 2017) and Piraux et al. (Piraux et al., 2018) traced QUIC’s progression over time. Additionally, Rüth et al. (Rüth et al., 2018) examined QUIC deployment and usage in its early stages, while QScanner (Zirngibl et al., 2021) analyzed early QUIC deployments.
@@
Features of QUIC Research on HTTP/2 and HTTP/3, such as the studies by Marx et al. (Marx et al., 2019) and Sander et al. (Sander et al., 2022), has provided in-depth analysis of how resource multiplexing and prioritization affect web performance. These studies highlight that while H3’s use of QUIC reduces HoL blocking compared to H2, real-world gains depend heavily on effective prioritization strategies. Other research has extended QUIC to multipath QUIC (MPQUIC) (Coninck and Bonaventure, 2017; Ingenzi et al., [n. d.]), and new packet schedulers for MPQUIC (do Prado Paiva et al., 2023; Rabitsch et al., 2018; Wang et al., 2019) and upper-layer solutions (Zheng et al., 2021) have followed. MultipathTester (Coninck and Bonaventure, 2019) compares the performance of MPTCP with MPQUIC on iOS. Several studies have focused on QUIC congestion control (Engelbart and Ott, 2021; Haile et al., 2022; Mishra et al., 2022), while others examine security and privacy (Kosek et al., 2022c; Lychev et al., 2015; Nawrocki et al., 2021, 2022; Soni and Rajput, 2020). Experimental protocols inspired by QUIC, like TCPLS (Rochet et al., 2021), DCQUIC (Tan et al., 2021), and MCQUIC (Navarre et al., 2023), further show its potential.
@@
QUIC with Proxy Exploring QUIC in proxy architectures is relatively sparse, but some studies have started to bridge this gap. Kosek et al. (Kosek et al., 2023) introduced Secure Middlebox-Assisted QUIC (SMAQ), enabling middleboxes in QUIC connections while preserving security, and demonstrated performance improvements in satellite networks. Another study by Kosek et al. (Kosek et al., 2022a) evaluated QUIC and HTTP/3 with and without performance-enhancing proxies (PEPs) in satellite networks, showing significant performance gains, especially in high-latency conditions. Kühlewind et al. (Kühlewind et al., 2021) assessed the QUIC-based MASQUE proxying protocol for tunneling UDP and IP traffic, revealing benefits like local recovery in lossy links despite introducing overhead. However, these studies focus on different aspects of QUIC and proxies. Our work uniquely compares H2 and H3 in proxy-enhanced environments, with a particular emphasis on QUIC’s connection migration feature, evaluating the performance impacts of proxy and connection migration under various network conditions.
@@
Single-Stream Request Proxygen (H3) Curl (H2, H3) Chrome (H2, H3) Wget (H2) Quiche (H3) aioquic (H3) Nginx Bandwidth: 10 Mbps Loss: 0, 2, 4 (%) Delay: 0, 25, 50 (ms) Small Files: 0KB, 10KB, 100KB Medium Files: 500KB, 1MB, 5MB Large Files: 10MB Total Request Time
@@
Web Page Loading Chrome (H2, H3) Nginx Bandwidth: 10 Mbps Loss: 0, 2, 4(%) Delay: 0, 25, 50 (ms) 18 popular local host websites Speed Index
3. Experimental Setup
We use a hardware network test bed that is remotely accessible for networking researchers and educators (Wiseman et al., 2008). The client and proxy machines run Ubuntu 20.04.6 LTS on Intel(R) Xeon(R) E5520 CPUs with 16GB RAM, while the server runs Ubuntu 22.04.4 LTS on an Intel(R) Core(TM) i3-3240 CPU with 16GB RAM. These machines are interconnected through a router via a 1 Gbps Ethernet network. The ’tc netem’ utility (tc, 2024) is used for bandwidth control, packet loss, and delay manipulation. In all experiments, the bandwidth is set to 10 Mbps, with specific loss and delay parameters detailed in each experiment’s setup (see Table 1). We use a fixed bandwidth to maintain controlled conditions, allowing us to isolate the effects of proxy environments on the performance of H2 and H3. These bandwidth limitations, combined with varied loss and delay parameters on both uplink and downlink, represent realistic edge-case network conditions that induce unique protocol behaviors for root cause analysis. The study by Zhang et al. (Zhang et al., 2024) has already explored performance under varying bandwidth, while both our work and studies like (Yu and Benson, 2021) use fixed bandwidth with varied loss and delay to focus specifically on how network impairments affect protocol performance, without introducing the additional complexity of fluctuating bandwidth.
@@ 3. Experimental Setup
In each experiment, we conducted 10 runs. We use Total Request Time as the metric in single-stream, multi-stream, and connection migration experiments, measuring from the client’s first request packet to the server until the client closes the connection. For web page loading, we use the Speed Index (spe, 2024), a metric that measures how quickly content appears on the user’s screen.
@@ 3.1. Baseline Experiment Setup
This experiment setup establishes the baseline for evaluating the performance of H2 and H3 in single-stream request, multi-stream request, connection migration, and web page loading. We employ Cubic as the CCA on the client side, and New Reno on the server side for both QUIC and TCP. Cubic is the default CCA for most H3 clients and the default CCA on Linux. For QUIC, the New Reno is the only CCA that Nginx supports. Both H2 and H3 use TLS 1.3. Our experimental setup is detailed in Table 1, with the logical topology illustrated in Figure 1.
Figure 1. Baseline Experiment Logical Topology
[…]
This figure shows the baseline experiment topology used in the study, illustrating the network structure and key components involved.
@@
[…]
This figure illustrates the sizes of different websites and the corresponding number of resources required to load them.
In the connection migration experiment, We extended Cloudflare’s Rust-based quiche client to support connection migration. Specifically, we took inspiration from the design of Andriod’s Connection Manager (Con, 2024) to detect changed interfaces/IP addresses, and initiate the connection migration process as described in (Iyengar and Thomson, 2022). For H2, we employed wget with and without range Requests (ran, 2024), along with curl tests without range requests. Range requests in H2 facilitate fetching a range of data. This allows us to explore H2’s ability to pause and resume downloading in the change in connection scenario. The experiment simulates IP address change after downloading over 2MB of a 5MB test file to mimic real-world network transitions, evaluating how each protocol maintains connection stability and download continuity during these transitions.
@@ 3.2. Proxy Experiment Setup
This experiment setup assesses the impact of introducing an Envoy reverse proxy between the server and client, with loss and delay applied between the proxy and the client. The reverse proxy breaks the connection between the client and the server and establishes a new connection with the server. We use BBR as the CCA for both QUIC and TCP since it is the default CCA for QUIC on Envoy. We disabled the proxy cache and retry features. The connection bandwidth is uniformly set at 10 Mbps across all links. We chose Envoy for its robust, production-grade capabilities and widespread adoption by major companies like Google, Microsoft, and Amazon. The network logical topology positions the Envoy between the client and server, as detailed in the accompanying Figure 3.
Figure 3. Proxy Experiment Logical Topology
[…]
This figure illustrates the topology used for the proxy experiment, showing the network structure and key components involved in the proxy-enhanced setup.
@@
Figure 4. Single-Stream Results. The times represent the mean values over 10 runs.
[…]
Single-Stream Results
@@ 4. Experimental Evaluation
In this section, we present the results of our experiments and our analysis of the results.
4.1. Single-Stream Experiment Evaluation
Figure 4(c) presents the results of our single-stream experiments across various network conditions, highlighting the performance differences between H2 and H3.
@@ 4.1. Single-Stream Experiment Evaluation
Moderate Network Impairments In scenarios involving moderate impairments (10 Mbps bandwidth, 0% packet loss, and 50 ms delay), both H2 and H3 demonstrated comparable performance without the use of a proxy (Figure 4(b)). However, The introduction of a proxy caused slight variations in performance for H3 clients, such as curl and quiche, but overall, H3 maintained consistent performance with minimal impact from the proxy. In contrast, H2’s performance remained stable across different clients, suggesting that H2 implementations are more mature and exhibit fewer discrepancies between implementations. In scenarios with loss only (10 Mbps bandwidth, 2-4% packet loss, 0 ms delay), both H2 and H3 exhibited comparable performance.
@@
Connection Establishment The QUIC connection starts with a handshake that merges the TCP 3-way handshake and the TLS 1.3 handshake into a single three-packet exchange. This process eliminates a full Round-Trip Time (RTT) in the QUIC startup phase. While both protocols show similar connection establishment times under ideal network conditions, H3’s advantages become evident under impaired conditions (Figure 4(c)). The quicker connection establishment does not significantly impact performance in optimal conditions, but it becomes crucial in severely impaired networks. For instance, under conditions of 4% packet loss and 50ms delay, H3 clients, on average, complete a 0KB file download 14.5% faster than their H2 counterparts.
@@
Impact of CCAs In proxy scenarios, the use of BBR can maintain high throughput and low latency by dynamically adjusting the congestion window based on real-time bandwidth and RTT estimations. This proactive congestion management is more effective than loss-based algorithms like Cubic, which only reduce the congestion window upon detecting packet loss. While BBR improves H2’s performance significantly in proxy-enhanced environments, this improvement is highly dependent on the correct CCA choice. In contrast, H3’s performance remains robust across different network conditions and CCAs, highlighting its superior consistency without relying on external configurations like BBR.
@@
Figure 5. Multi-Stream Results. Each point represents the highest Total Request Time among 20 test files for each run. A total of 10 runs were conducted for each loss level.
[…]
Multi-Stream Results.